OpenStudy (anonymous):

Need help with this expanded notation question, would the number 2,435 be 2(1000)+4(100)+3(10)+5(1)= 2,435? Please help. Thank you!

OpenStudy (mertsj):

yes

OpenStudy (lilg132):

2 * 1000 = ? 4 * 100 = ? 3 * 10 = ? 5 * 1 = ?

OpenStudy (lilg132):

that is how you expand the brackets and check